# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v1

@@ V-B2 Compared with weighted cross-entropy loss
As depicted in Table IV, AUNCE has exhibited superior performance compared to the traditional weighted cross-entropy loss (WCE) for AU detection, resulting in an increase in F1-score from 63.6% to 65.7% on BP4D and from 62.2% to 66.4% on DISFA. This highlights that capturing the distinctive features of different samples proves more effective than concentrating on pixel-level information. Consequently, our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance problem for each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection.
AUNCE can seamlessly integrate into various lightweight AU detection networks, serving as a replacement for generative models. The loss demonstrates robust compatibility in diverse model architectures.
V-C Hyperparameters Analysis
@@
TABLE V: Quantitative results the probabilities of various positive sample pairs $p_{1},p_{2},p_{3},p_{4}$ on BP4D dataset.
$p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ F1-Score
@@
V-C1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
@@
The experimental results are presented in Table V. When $p_{1}$=$p_{2}$=0.2, $p_{3}$=0.6, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent errors and noise. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of error and noisy samples, thereby reinforcing the robustness of the training process.
Refer to caption (a)
Refer to caption (b)
Figure 3: Quantitative results of $\beta$ of different positive and negative samples for each AU type on BP4D dataset. From up to down: (a) $\beta$ for positive samples (the minority class), and (b) $\beta$ for negative samples (the majority class).
V-C2 Backpropagation rate controller $\beta$
As is well understood, the hyperparameter $\beta$ plays a crucial role in governing the relative importance of positive and negative samples during the parameter update process. It ensures that the model allocates more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. To explore its effects, we systematically vary the hyperparameter $\beta$ for the minority class from 0.8 to 1.8, with $\beta$ for the majority class fixed at 0.4. Additionally, we vary $\beta$ for the majority class from 0.2 to 1.2, keeping $\beta$ for the minority class fixed at 1.2.
The experimental results are depicted in Fig. 3. The optimal experimental outcome is observed when $\beta$ for the minority class is set to 1.2, and $\beta$ for the majority class is fixed at 0.4. Hence, when $\beta_{\text{minority}}>\beta_{\text{majority}}$, the impact of the imbalanced problem can be effectively mitigated.
VI Conclusion
In this paper, we introduce an innovative discriminative contrastive learning paradigm designed to replace traditional pixel-level learning in the domain of AU detection. Our proposed contrastive loss, referred to as AUNCE, shares similarities with InfoNCE, transforming the task of learning image data distribution information into the more effective learning of differentiated information among images. Additionally, we present a positive sample sampling strategy that introduces three types of positive sample pairs, addressing the challenge of AU labels tainted by noise and errors in two AU datasets. To further enhance the framework, we incorporate a importance re-weighting strategy specifically tailored to minority AUs. This strategy alters the backpropagation gradient of positive and negative samples, effectively addressing the imbalanced nature of each AU type. Experimental results underscore the superior performance of AUNCE in comparison to existing methods.
@@ Acknowledgement
The computation is completed in the HPC Platform of Huazhong University of Science and Technology.

### v3

@@ IV-C2 Evaluating other components
Based on all ablation studies, it demonstrates that our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance issue of each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection. The introduction of each component proves beneficial for performance improvement, with NS emerging as the most effective component. Moreover, the results on each AU are also shown in Table VII. It can be seen that the results on basically every AU have improved or approached after we add $w_{i}$, PS and NS.
IV-D Hyperparameters Analysis
@@
TABLE VIII: Quantitative results the probabilities of various positive sample pairs $p_{1},p_{2},p_{3},p_{4}$ on BP4D dataset.
$p_{1}$ $p_{2}$ $p_{3}$ $p_{4}$ F1-Score(%)
@@
IV-D1 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
@@
The experimental results are presented in Table VIII. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The conclusion is consistent with  [89, 90]. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
Figure 5: Quantitative results of $\beta$ of different positive and negative samples for each AU type on BP4D dataset. From left to right: (a) $\beta$ for positive samples (the minority class), and (b) $\beta$ for negative samples (the majority class).
IV-D2 Backpropagation rate controller $\beta$
As is well understood, the hyperparameter $\beta$ plays a crucial role in governing the relative importance of minority and majority class samples during the parameter update process. It ensures that the model allocates more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. To explore its effects, we systematically vary the hyperparameter $\beta$ for the minority class from 0.8 to 1.8, with $\beta$ for the majority class fixed at 0.4. Additionally, we vary $\beta$ for the majority class from 0.2 to 1.2, keeping $\beta$ for the minority class fixed at 1.2.
The experimental results are depicted in Fig. 5. The optimal experimental outcome is observed when $\beta$ for the minority class is set to 1.2, and $\beta$ for the majority class is fixed at 0.4. Hence, when $\beta_{\text{minority}}>\beta_{\text{majority}}$, the impact of the class imbalance issue can be effectively mitigated.
V Conclusion
In this paper, we introduce a contrastive learning framework for AU detection. Our proposed contrastive loss, named AUNCE, shares similarities with InfoNCE, transforming the task of learning entire pixel-level information of AUs into the more effective learning of subtle feature differences when AUs activate. Additionally, we present a positive sample sampling strategy that introduces three types of positive sample pairs, addressing the challenge of AU labels tainted by noise and errors in two AU datasets. To further enhance the framework, we incorporate a negative sample re-weighting strategy specifically aimed for minority AUs. This strategy controls the relative contribution of minority class samples to the parameter updates, effectively addressing the class imbalance issue of each AU type, and contributing to hard negative sample mining. Experimental results underscore the superior performance of AUNCE in comparison to existing methods.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Science and Technology Program (Nos. 2024NSFTD0036, 2024ZHCG0166). The computation is completed in the HPC Platform of Huazhong University of Science and Technology.